# Monitoring of water volume in a porous reservoir using seismic data: Validation of a numerical model with a field experiment

### v3

@@ Abstract
As global groundwater levels continue to decline rapidly, there is a growing need for advanced techniques to monitor and manage aquifers effectively. This study focuses on validating a numerical model using seismic data from a small-scale experimental setup designed to estimate water volume in a porous reservoir. Expanding on previous work with synthetic data, we analyze seismic data acquired from a controlled experimental site in Laukaa, Finland. By employing neural networks, we directly estimate water volume from seismic responses, bypassing the traditional need for separate determinations, for example, of reservoir water table level and porosity. The study models wave propagation through a coupled poroviscoelastic-viscoelastic medium using a three-dimensional discontinuous Galerkin method. The proposed methodology is validated against experimental data, aiming to improve precision in mapping current water volumes and contributing to the development of sustainable groundwater management practices.
@@ 1 Introduction
Groundwater aquifers are facing unprecedented threats, with levels decreasing at alarming rates, often more than one meter per year in some areas—leading to significant long-term reductions. As a result, surface water flows that were previously sustained by groundwater are becoming seasonal or disappearing completely [9]. To ensure sustainable water extraction, it is crucial to have a better understanding of the location and extent of groundwater resources. In response to these challenges, this study focuses on validating a numerical model using seismic data from a small-scale experimental setup with emphasis to estimate water volume in a porous reservoir at various water table levels. Neural networks have been increasingly applied to enhance the analysis of seismic data, for a recent review, see [19]. By using neural network techniques, we aim to directly estimate water volume from seismic responses, avoiding the traditional need for separate determinations, for example, of reservoir water table level and porosity.
@@
Building upon previous works with synthetic data in both two [18] and three spatial dimensions [14], this study focuses on the estimation of water volume from seismic data collected at a controlled experimental site in Laukaa, Finland. The seismic data, obtained from an artificial porous sand pool using a drop-weight seismic source, were gathered during several acquisition campaigns following changes in the water table level. For the neural network-based water volume estimator, we first build a synthetic training database by simulating seismic wave propagation for different scenarios of the studied sand pool using a numerical wave propagation solver. The applied numerical wave propagation solver builds upon the works presented in [5, 14]. This synthetic data is then used to train neural networks, which are subsequently applied to the real seismic data to directly recover water volume, aiming to improve precision in mapping current water volumes and contributing to the development of sustainable groundwater management practices.
@@
In our analysis of neural network-based estimates, we augment our approach by applying the Shapley additive explanation (SHAP) framework [17]. Employing the SHAP framework provides us with a deeper understanding of the estimation process at the receiver level. In our specific context, we focus on understanding the contribution of receivers to our results, rather than aiming to optimize the configuration of the receiver array.
We present a comprehensive study on employing neural networks to characterize water storage using seismic data from the Laukaa test site. Section 2 describes the field measurements at the Laukaa site. In Section 3, we discuss the simulation of measurement data and the synthetic modeling of the site. The use of neural networks to characterize water storage is discussed in Section 4. Section 5 offers a thorough analysis of the results, with a focus on the water volume predictions and the results of the SHAP analysis. In Section 6, we summarize our findings and suggest possible directions for future research.
2 Laukaa test site - field measurements
2.1 Description of the Laukaa test site
The field measurements are conducted in a man-made sand pool located at Natural Resources Institute Finland (Luke) premises in Laukaa, Finland. Generally, sand pool serves a controlled environment to explore groundwater distribution with a knowledge of the media’s geometry and physical parameters. The Laukaa test site stands as a homogeneous and isotropic custom pool, integrated into a recirculating aquaculture system. It contains uniform sand grain size and is surrounded by an impermeable clay lining, as detailed in [21]. The pool’s dimensions are well-defined, facilitating adjustments to the groundwater table level as required. With this setup, it is possible to calculate the actual volumes of water as well as collect the seismic data corresponding to them.
2.2 Seismic measurements
The seismic measurements were conducted in Laukaa in June 2022. The measurements from the gauge shallow wells were collected to get the ground truth values to calculate the water volume. To study different water levels, we opened a valve to reduce the water table level, followed by a period of stabilization. The stabilization process was monitored by measuring the water table levels from different wells placed along the groundwater flow path.
@@ 2.2 Seismic measurements
In our experiments, a metallic rod weight was dropped on a steel plate and served as a source of seismic waves, with an electric brake mechanism to control the rod’s release and prevent multiple hits. The experiments involved shots at 13 specific locations and were conducted at three different drop heights ($\rm{dh}$), consisting of 5, 10, and 15 cm elevations. Three different heights were selected to test the signal amplitude clipping threshold at the nearest seismic receivers.
For data acquisition, three-component (3C) 5 Hz geophones connected to commercially available 24-bit nodal seismic recorders were used to collect the data. The recorders were deployed along 4 receiver lines (14 per line) with 0.5 m inline and 1.5 m crossline spacing, plus an extra receiver randomly positioned, making a rather fine-scale 3D seismic setup. The recorded time length was 3 seconds with 4 kHz sampling rate.
@@
Figure 1: Photo of experimental setup at Laukaa test site, showing the grid of geophones, weight-drop source, and wells to measure the water table level.
Various sources of noise were active throughout the measurements, such as river currents and ongoing construction work contributing to the background noise. Additionally, the occasional operation of lawnmowers was also present on the site. To minimize potential interference and enhance data quality, the measurements were deliberately scheduled during rain-free days.
Figure 1 displays an image of the Laukaa test site, illustrating the arrangement of geophones, the weight-drop source tool, and wells used to gauge the water table level (black pipes). The receiver setup includes a total of 57 3C geophones (blue), each accompanied by a nodal seismic recorder (white box).
@@ 3 Modeling of the test site and field measurements
The model’s geometry for synthetic data generation replicates the Laukaa sand pool, capturing its structure and physical parameters. This replicated model serves as a tool for generating extensive synthetic training data for our neural network model. To generate synthetic seismograms, we employ Biot’s isotropic poroviscoelastic model to handle wave propagation in the porous medium. In a zone adjacent to the porous material, the isotropic viscoelastic model is utilized.
@@ 3.1 Synthetic model of the Laukaa test site
The applied synthetic model shown in Fig. 2 replicates the geometry of the Laukaa test site. It is a box with a length of 31.5 m, a width of 16.2 m, and a height of 2.75 m. In addition, the corners of the geometry are rounded as shown in the graph. The maximum length and width of the air-saturated zone are 29.5 m and 14.2 m, respectively. Finally, the bottom profile of the water-saturated zone is a rectangle with a length of 23.5 m and a width of 8.2 m. The bottom is located at a depth of 2 m from the top surface. A water table divides the porous material into air-saturated and water-saturated subdomains. When creating the data with the synthetic model, the water table level $z$ is randomized from $z\sim\mathcal{U}(-120,-25)$ cm.
@@
Figure 2: A schematic of the problem geometry with the top showing cross-section and the bottom an oblique angle 3D view. Light-blue color refers to air-saturated and dark-blue to water-saturated zone. Light-brown color denotes the surrounding elastic material. The setup contains a total of 57 receivers, marked with red dots and one green dot, all of which are located on the ground surface. The green dot serves as the reference receiver and is located between seventh receiver on lines 1 and 2. The source location is marked with a yellow star and placed close to the fourth receiver on line 1.
@@
Our model includes a total of 57 receivers that record the solid velocity components in the horizontal ($y$) and vertical ($z$) directions, denoted by $v_{\rm s}$ and $w_{\rm s}$, respectively. These components represent the solid velocity in the model’s coordinate system (see Fig. 2). Notably, they do not correspond to isolated wave modes (e.g., P-, SV-, or SH-waves), but rather capture the full wavefield projected onto the model axes. As a result, the signals may contain contributions from multiple wave types, due to mode conversions and the complex nature of wave propagation in porous media.
Exact receiver locations were accurately surveyed using a commercially available differential GPS (DGPS) system and are used in the numerical modeling. One additional receiver, marked in green, is located near line 1 and serves as a reference point. A more detailed discussion of this reference point is provided in Section 4.1.
@@
Although the measurement data acquisition involved 13 shot locations, the water volume estimation is based on data from a single source location only, as used in the simulation study [14]. The location for the source was arbitrarily selected to be the closest to receiver 4 on Line 1. The seismic source is modeled as a vertical force pointing to the negative $z$-axis. Since the exact location of the source slightly varies between different measurements (distance from the closest geophone was measured with a ruler), the location was assumed to be uncertain in the numerical model when training the neural network. In practice, to account for uncertainty in repositioning the source between measurements at different water table levels, we assume a 10 cm variability in the source location in both the $x$ and $y$ directions. This means that the source center location in the $(x,y)$-plane is randomized as $x\sim\mathcal{U}(4.4890,4.5890)$ m and $y\sim\mathcal{U}(-5.2142,-5.1142)$ m. Receivers and the source are placed on the ground surface.
As a source function, we use a first derivative of a Gaussian
$g=\frac{(t-t_{0})}{c}\exp\left(b\left(\left(t-t_{0}\right)^{2}-c^{2}\right)\right),$ (1)
where $b=-(f\pi)^{2}$ and $c=\sqrt{-0.5/b}$. When creating the training data, we set frequency $f$ to 60 Hz, time delay $t_{0}$ to $1.2/f$, and modelling time to 0.35 s.
@@ 3.2 Physical parameters
The parameters defining the viscoelastic material are based on solid density $\rho_{\mathrm{e}}$, pressure wave speed $c_{P}$, shear wave speed $c_{S}$, and quality factors $Q_{P}$ and $Q_{S}$. The quality factors define the level of viscous attenuation in the medium. In the current paper, these parameters are randomized from uniform distributions, and minimum and maximum values are given in Table 1. One must note, that the attenuation is modeled with three mechanisms in both porous reservoir and the surrounding medium, see [14].
Table 1: Material parameter bounds assumed for the elastic material.
variable name symbol (unit) Minimum value Maximum value
@@
Quality factor $Q_{P}$ 20 50
Quality factor $Q_{S}$ 20 50
[…]
Permeability $k$ is calculated from the Darcy law
$\frac{k}{\eta}=\frac{K}{\rho_{f}g},$ (2)
where $g=9.81$ m s-2. In this work, the hydraulic conductivity $K$ is approximated from [3]
$K\,(\mbox{m/s})=a\left(I_{0}+0.025\left(d_{50}-d_{10}\right)\right)^{2},$ (3)
where $a=1300/(24\cdot 60^{2})$ and $I_{0}$ is the intercept of the line formed by grain-size values with the grain-size axis. Minimum and maximum values for the grain-size parameters $d_{10}$ and $d_{50}$ are given in Table 2. We compute the permeability value for each material sample using the viscosity and density values for water.
Table 2: Material parameter bounds assumed for the water reservoir.
variable name symbol (unit) Minimum value Maximum value
Mass density of sand grains $\rho_{\mathrm{s}}$ (kg m-3) 2400 2800
@@
Quality factor $Q_{\mu_{\mathrm{fr}}}$ 15 50
Quality factor $Q_{\kappa_{\mathrm{s}}}$ 80 120
Quality factor $Q_{\kappa_{\mathrm{fr}}}$ 15 50
Grain-size $d_{10}$ (mm) 0.4 0.8
Grain-size $d_{50}$ (mm) 1.1 1.6
@@
For the Laukaa test case studied in this paper, the parameter distributions shown in Tables 1 and 2 are chosen to be representative of the geological properties observed in the region. However, it is important to note that while these ranges are designed to capture key variations, the neural network model may still perform well even if not all parameters are completely accounted for in the training data. This is because the network is able to learn relevant features from the available data, potentially allowing it to generalize and provide accurate estimates even in regions not explicitly covered by the training set. In this sense, the model can adapt to variations in parameters and still maintain robustness in its predictions [10].
3.3 Computing of seismic data
In this research, we utilize an in-house software that is based on the discontinuous Galerkin (DG) method [12] and the third-order Adams-Bashforth time-stepping [6] techniques to generate synthetic seismograms. The applied software employs tetrahedral elements to discretize the geometry of the problem. In addition to the element size, the accuracy can be controlled by selecting the order of the polynomial basis functions. For an in-depth discussion of the applied software and the associated methodology, we refer to [14] and references therein.

### v4

@@ Abstract
As global groundwater levels continue to decline rapidly, there is a growing need for advanced techniques to monitor and manage aquifers effectively. This study focuses on validating a numerical model using seismic data from a small-scale experimental setup designed to estimate water volume in a porous reservoir. Expanding on previous work with synthetic data, we analyze seismic data acquired from a controlled experimental site in Laukaa, Finland. By employing neural networks, we directly estimate water volume from seismic responses, bypassing the traditional need for separate determinations, for example, of reservoir water-table level and porosity. The study models wave propagation through a coupled poroviscoelastic-viscoelastic medium using a three-dimensional discontinuous Galerkin method. The proposed methodology is validated against experimental data, aiming to improve precision in mapping current water volumes and contributing to the development of sustainable groundwater management practices.
@@ 1 Introduction
Groundwater aquifers are facing unprecedented threats, with levels decreasing at alarming rates, often more than one meter per year in some areas—leading to significant long-term reductions. As a result, surface water flows that were previously sustained by groundwater are becoming seasonal or disappearing completely [10]. To ensure sustainable water extraction, it is crucial to have a better understanding of the location and extent of groundwater resources. Motivated by these challenges, this study explores the potential to estimate water volume in an artificial porous reservoir from seismic measurements using a combination of numerical modeling and neural networks, and validates the approach using controlled field data. Neural networks have been increasingly applied to enhance the analysis of seismic data, for a recent review, see [26]. By using neural network techniques, we aim to directly estimate water volume from seismic responses, avoiding the traditional need for separate determinations, for example, of reservoir water-table level and porosity.
@@
Building upon previous works with synthetic data in both two (2D) [19] and three (3D) spatial dimensions [17], this study focuses on the estimation of water volume from seismic data collected at a controlled experimental site in Laukaa, Finland. The seismic data, obtained from an artificial porous sand pool using a drop-weight seismic source, were gathered during several acquisition campaigns following controlled changes in the water-table level. For the neural network-based water volume estimator, we first build a synthetic training database by simulating seismic wave propagation for different scenarios of the studied sand pool, including variations in physical parameters and water table level. The simulations employ a numerical wave propagation solver based on the methods presented in [7, 6, 17]. This synthetic data is then used to train neural networks, which are subsequently applied to the real seismic data to directly recover water volume, aiming to improve precision in mapping current water volumes and contributing to the development of sustainable groundwater management practices.
@@
In our analysis of neural network-based estimates, we augment our approach by applying the Shapley additive explanation (SHAP) framework [22]. Employing the SHAP framework provides us with a deeper understanding of the estimation process at the receiver level. Specifically, we use SHAP analysis to evaluate the contribution of each receiver to the final water-volume estimate. The purpose is interpretability rather than receiver-array optimization.
In this study, we present a comprehensive framework for estimating water volume in a porous reservoir using seismic data and neural networks, validated through field experiments at the aforementioned Laukaa test site. Section 2 provides an overview of the site and details the seismic measurements conducted under varying water table conditions. In Section 3, we describe the construction of a 3D numerical model that replicates the test site, used for simulating synthetic seismograms based on Biot’s poroviscoelastic theory. Section 4 introduces the neural network architecture, including the data preprocessing steps, training methodology, and interpretability analysis using SHAP values. Results are shown in Section 5, with a focus on the water volume predictions and the SHAP analysis. In Section 6, we summarize our findings.
2 Site description and seismic measurements
The field measurements were conducted in a man-made sand pool at the Natural Resources Institute Finland (Luke) in Laukaa, Finland. Generally, the sand pool serves as a controlled environment to explore groundwater distribution with a knowledge of the media’s geometry and physical parameters. Although natural small-scale variability exists, the site can be reasonably approximated as a homogeneous and isotropic sand body for the purposes of our modelling. For conceptual clarity, we divide the domain into three approximate zones: an air-saturated porous region, a water-saturated porous region, and the surrounding elastic material. The pool is enclosed by an impermeable clay lining [28] and its dimensions are well defined, allowing controlled adjustments of the water-table level. This makes it possible to compute the true water volumes for each measurement configuration. The water-saturated region occupies the bottom of the pool (23.5 m by 8.2 m at 2 m depth), while the overlying air-saturated zone extends to approximately 29.5 m by 14.2 m.
@@ 2 Site description and seismic measurements
The seismic measurements were conducted in June 2022. In our experiments, a metallic rod weight was dropped on a steel plate and served as a source of seismic waves, with an electric brake mechanism to control the rod’s release and prevent multiple hits. The experiments involved shots at 13 specific locations and were conducted at three different drop heights ($\rm{dh}$): 5, 10, and 15 cm. The different drop heights were used to vary the source amplitude and check for potential signal clipping at the nearest receivers.
For data acquisition, three-component (3C) 5 Hz geophones connected to 24-bit nodal seismic recorders were used, obtained from the Finnish Seismic Instrument Pool (FINNSIP) [14]. Various sources of ambient noise were present during the field campaign, including river currents, nearby construction activity, and occasional lawnmower operation. To reduce their influence, measurements were scheduled during rain-free and relatively quiet periods. The instruments were deployed along four receiver lines (14 receivers per line) with approximately 1.0 m inline and 1.7 m crossline spacing, plus one additional station placed at a random location, forming a fine-scale 3D acquisition layout. Exact receiver locations were surveyed using Differential Global Positioning System (DGPS) and incorporated directly into the numerical simulations. The seismic data were recorded at 4 kHz sampling rate.
@@
Figure 1: Photo of experimental setup at Laukaa test site, showing the grid of geophones, weight-drop source, and wells to measure the water-table level.
Figure 1 shows the Laukaa test site, including the layout of the geophones, the weight-drop source, and the black pipes housing the water-table gauge wells. A total of 57 three-component geophones (blue) and their corresponding nodal recorders (white boxes) were deployed. To obtain ground-truth water-volume estimates, water-table levels were monitored using these gauge wells at several locations around the pool. The water level was adjusted by releasing water to the adjacent river, after which the system was allowed to stabilize before seismic acquisition. These auxiliary measurements were used exclusively for validating the estimated water volumes.
@@ 3 Modeling of the test site and field measurements
This section presents a three-dimensional (3D) synthetic model developed to simulate wave propagation and generate training data for the neural network. The model’s geometry for synthetic data generation replicates the Laukaa sand pool, capturing its structure and physical parameters. This replicated model serves as a tool for generating extensive synthetic training data for our neural network model. To generate synthetic seismograms, we employ Biot’s isotropic poroviscoelastic model to handle wave propagation in the porous medium. In a zone adjacent to the porous material, the isotropic viscoelastic model is utilized.
3.1 Governing equations
We model wave propagation in porous and elastic media using well-established partial differential equations (PDEs). Here, we summarize the main PDE systems and the associated physical parameters. More detailed derivations, numerical implementations, and attenuation models are available in [17] and references therein.
3.1.1 Biot’s poroelastic wave equation
The coupled dynamics of solid and fluid phases in porous media are described by Biot’s equations [3]
$\rho_{\mathrm{a}}\frac{\partial^{2}\mathbf{u}_{\rm s}}{\partial t^{2}}+\rho_{\mathrm{f}}\frac{\partial^{2}\mathbf{w}}{\partial t^{2}}$ $=\nabla\cdot\mathbf{T},$ (1)
$\rho_{\mathrm{f}}\frac{\partial^{2}\mathbf{u}_{\rm s}}{\partial t^{2}}+m\frac{\partial^{2}\mathbf{w}}{\partial t^{2}}+\frac{\eta}{k}\frac{\partial\mathbf{w}}{\partial t}$ $=\nabla\cdot\mathbf{T}_{\rm f},$ (2)
where $\mathbf{u}_{\rm s}$ is the solid displacement, and $\mathbf{w}=\phi(\mathbf{u}_{\rm f}-\mathbf{u}_{\rm s})$ is the relative fluid displacement (scaled by porosity $\phi$). Here, $\mathbf{u}_{\rm f}$ denotes the fluid displacement. The average density is $\rho_{\mathrm{a}}=(1-\phi)\rho_{\mathrm{s}}+\phi\rho_{\mathrm{f}}$, where $\rho_{\mathrm{f}}$ is the fluid density, $\rho_{\mathrm{s}}$ is the solid density, $\eta$ is the fluid viscosity, and $k$ is the permeability. The parameter $m=\rho_{\mathrm{f}}\tau/\phi$, with tortuosity $\tau$.
The total stress tensor $\mathbf{T}$ and fluid stress tensor $\mathbf{T}_{\rm f}$ are given by
$\mathbf{T}$ $=2\mu_{\mathrm{fr}}\mathbf{E}+\big(\lambda\operatorname{tr}(\mathbf{E})-\alpha_{\rm B}M\zeta\big)\mathbf{I},$ (3)
$\mathbf{T}_{\rm f}$ $=\big(\alpha_{\rm B}M\operatorname{tr}(\mathbf{E})-M\zeta\big)\mathbf{I},$ (4)
where $\mathbf{E}$ is the solid strain tensor and $\zeta=-\nabla\cdot\mathbf{w}$ is the fluid content variation. The parameter $\mu_{\mathrm{fr}}$ is the shear modulus of the frame, and $\mathbf{I}$ denotes the identity tensor.
The remaining parameters, $\lambda$, $\alpha_{\rm B}$, and $M$, are defined as
$\lambda$ $=\kappa_{\mathrm{fr}}+\alpha_{\rm B}^{2}M-\frac{2}{3}\mu_{\mathrm{fr}},$ (5)
$\alpha_{\rm B}$ $=1-\frac{\kappa_{\mathrm{fr}}}{\kappa_{\mathrm{s}}},$ (6)
$M$ $=\frac{\kappa_{\mathrm{s}}}{\alpha_{\rm B}-\phi(1-\kappa_{\mathrm{s}}/\kappa_{\mathrm{f}})},$ (7)
where $\kappa_{\mathrm{fr}}$, $\kappa_{\mathrm{s}}$, and $\kappa_{\mathrm{f}}$ are the frame, solid, and fluid bulk moduli, respectively.
3.1.2 Elastic wave equation
For purely elastic media, the displacement $\mathbf{u}_{\rm e}$ satisfies
$\rho_{\mathrm{e}}\frac{\partial^{2}\mathbf{u}_{\rm e}}{\partial t^{2}}=\nabla\cdot\mathbf{S},$ (8)
with stress tensor
$\mathbf{S}=2\mu_{\mathrm{e}}\mathbf{E}+\lambda_{\mathrm{e}}\operatorname{tr}(\mathbf{E})\mathbf{I},$ (9)
where $\rho_{\mathrm{e}}$ is the density, and $\mu_{\mathrm{e}}$, $\lambda_{\mathrm{e}}$ are the Lamé parameters. These parameters determine the compressional and shear wave speeds
$c_{P}=\sqrt{\frac{\lambda_{\mathrm{e}}+2\mu_{\mathrm{e}}}{\rho_{\mathrm{e}}}},\quad c_{S}=\sqrt{\frac{\mu_{\mathrm{e}}}{\rho_{\mathrm{e}}}}.$
At interfaces between elastic and poroelastic media, we follow the coupling strategy used in [7, 6]. The interface states are obtained by solving the corresponding elastic–poroelastic Riemann problem, subject to the physically required continuity conditions: continuity of solid velocity (normal and tangential) and continuity of total traction, while the normal fluid velocity vanishes on the elastic side.
3.1.3 Modeling attenuation
Attenuation in elastic and poroelastic media arises from different physical mechanisms. In both cases, stiffness attenuation is incorporated using a generalized Maxwell body (GMB) rheology [25]. The GMB introduces additional free parameters, specifically relaxation frequencies and memory variables, which govern the temporal behavior of stress relaxation and allow frequency-dependent attenuation. In elastic media, these parameters are tuned to match the desired quality factors $Q_{P}$ and $Q_{S}$. In poroelastic media, the quality factors $Q_{\mu_{\mathrm{fr}}}$, $Q_{\kappa_{\mathrm{fr}}}$, $Q_{\kappa_{\mathrm{s}}}$, and $Q_{\kappa_{\mathrm{f}}}$ are included in the model to account for frequency-dependent attenuation of the frame, solid, and fluid components. The full formulation of the GMB-based attenuation model, including how these memory variables are coupled with the governing PDEs and how the unrelaxed material parameters are derived, is provided in [17] and is not repeated here. For completeness, we note that the attenuation formulation used in [17] follows the poroelastic and viscoelastic GMB developments presented in [33, 32].
In poroelastic media, viscodynamic attenuation also plays a role, arising from viscous drag between fluid and solid phases. In this work, we operate within Biot’s low-frequency regime, where the dominant dissipation mechanism is frequency-independent fluid flow resistance. This results in a stiff PDE system due to the large contrast between relaxation and wave propagation timescales. We express the governing equations in first-order form using a velocity-strain formulation, leading to a hyperbolic system with stiff source terms. To handle this numerically, we apply Godunov splitting [21] to separate the dissipative and conservative components. Following [30], we solve the stiff dissipative terms analytically at each time step, while integrating the remaining hyperbolic system using the wave propagation software described in Section 3.4.
@@ 3.2 Synthetic model of the Laukaa test site
The applied synthetic model represents a finite section of the porous medium used for numerical simulations. It is a rectangular box with a length of 31.5 m, a width of 16.2 m, and a height of 2.75 m, with slightly rounded corners as shown in Fig. 2. It includes realistic air- and water-saturated subdomains, and a viscoelastic surrounding. The water-table level $z$ in the synthetic model is randomized as $z\sim\mathcal{U}(-120,-25)$ cm from the top surface.
@@
Figure 2: A schematic of the problem geometry with the top showing cross-section and the bottom an oblique angle 3D view. Light-blue colour refers to air-saturated and dark-blue to water-saturated zone. Light-brown colour denotes the surrounding elastic material. The setup contains a total of 57 receivers, marked with red dots and one green dot, all of which are located on the ground surface. The green dot serves as the reference receiver and is located between the seventh receiver on lines 1 and 2. The source location is marked with a yellow star and placed close to the fourth receiver on line 1. This figure is motivated by the problem geometry presented in [17].
@@
Our model includes a total of 57 receivers that record the solid velocity components in the horizontal ($y$) and vertical ($z$) directions, denoted by $v_{\rm s}$ and $w_{\rm s}$, respectively. These components represent the solid velocity in the model’s coordinate system (see Fig. 2). Notably, they do not correspond to isolated wave modes (e.g., P-, SV-, or SH-waves), but rather capture the full wavefield projected onto the model axes. As a result, the signals may contain contributions from multiple wave types, due to mode conversions and the complex nature of wave propagation in porous media. One additional receiver, marked in green, is located between the lines 1 and 2 which serves as a reference point. A more detailed discussion of this reference point is provided in Section 4.1.
@@
Although the measurement data acquisition involved 13 shot locations, the water volume estimation is based on data from a single source location only, as used in the simulation study [17]. The location for the source was arbitrarily selected to be the closest to receiver 4 on Line 1. The seismic source is modeled as a vertical force pointing to the negative $z$-axis. Since the exact location of the source slightly varies between different measurements (distance from the closest geophone was measured with a ruler), the location was assumed to be uncertain in the numerical model when training the neural network. In practice, to account for uncertainty in repositioning the source between measurements at different water-table levels, we assume a 10 cm variability in the source location in both the $x$ and $y$ directions. This means that the source center location in the $(x,y)$-plane is randomized as $x\sim\mathcal{U}(4.4890,4.5890)$ m and $y\sim\mathcal{U}(-5.2142,-5.1142)$ m. Receivers and the source are placed on the ground surface.
@@ 3.3 Physical parameters
The computational domain consists of three layers: an elastic surrounding medium and two porous subdomains, which are water-saturated and air-saturated. The elastic medium is modeled as viscoelastic, with properties defined by solid density $\rho_{\mathrm{e}}$, pressure and shear wave speeds ($c_{P}$, $c_{S}$), and quality factors $Q_{P}$ and $Q_{S}$, which account for attenuation. The porous subdomains share the same solid matrix properties, including grain density, porosity, tortuosity, solid and frame moduli, and grain-size distribution. All of these material parameters are randomized from uniform distributions, with bounds provided in Table 1. The porous reservoir parameter values in Table 1 are based on site-specific information provided in the internal report (in Finnish and not publicly available) of the Geological Survey of Finland (GTK) as well as [28, 23]. The surrounding material (marked with light-brown in Fig. 2) is approximated by an idealized elastic medium whose parameters are selected based on the original structural descriptions of the site.
The distinction between the two porous layers lies in the fluid phase: the water-saturated domain uses fixed values for fluid density, bulk modulus, and viscosity specific to water, while the air-saturated domain uses different fixed values representative of air. These fluid parameters are also listed in Table 1.
Permeability $k$ is derived from the Darcy law
$\frac{k}{\eta}=\frac{K}{\rho_{f}g},$ (10)
where $g=9.81$ m s-2, and $K$ is the hydraulic conductivity estimated from grain-size parameters $d_{10}$ and $d_{50}$ following [2]
$K\,(\mbox{m/s})=a\left(I_{0}+0.025(d_{50}-d_{10})\right)^{2},$ (11)
with $a=1.505\times 10^{-3}$, and $I_{0}$ denoting the extrapolated lower-limit grain size [2]. The grain-size parameters are also randomized as shown in Table 1. We compute the permeability value for each material sample using the viscosity and density values for water. One must note, that the attenuation is modeled with three mechanisms in both porous reservoir and the surrounding medium, see [17].
Table 1: Material parameters used in the three-layer model. Layers: (1) elastic surrounding medium, (2) porous reservoir (water-saturated), (3) porous zone (air-saturated). Porous layers share the same solid matrix properties.
Layer Parameter Symbol (unit) Min Max
Layer 1: Elastic surrounding medium (viscoelastic)
@@
Quality factor (P-wave) $Q_{P}$ 20 50
Quality factor (S-wave) $Q_{S}$ 20 50
Layers 2–3: Porous matrix (shared by both saturated zones)
Grain density $\rho_{\mathrm{s}}$ (kg m-3) 2400 2800
@@
Quality factor (Shear modulus) $Q_{\mu_{\mathrm{fr}}}$ 15 50
Quality factor (Solid bulk modulus) $Q_{\kappa_{\mathrm{s}}}$ 80 120
Quality factor (Frame bulk modulus) $Q_{\kappa_{\mathrm{fr}}}$ 15 50
Quality factor (Fluid bulk modulus) $Q_{\kappa_{\mathrm{f}}}$ $\infty$
Grain size $d_{10}$ ($m$) 0.4 0.8
Grain size $d_{50}$ ($m$) 1.1 1.6
Layer 2 only: Fluid parameters (water-saturated)
Fluid density $\rho_{\mathrm{f}}$ (kg m-3) 1000
Fluid bulk modulus $\kappa_{\mathrm{f}}$ (GPa) 2.1025
@@
Layer 3 only: Fluid parameters (air-saturated)
Fluid density $\rho_{\mathrm{f}}$ (kg m-3) 1.2
Fluid bulk modulus $\kappa_{\mathrm{f}}$ (MPa) 0.13628
[…]
@@
For the Laukaa test case studied in this paper, the parameter distributions shown in Table 1 are chosen to be representative of the geological properties observed in the region. However, it is important to note that while these ranges are designed to capture key variations, the neural network model may still perform well even if not all parameters are completely accounted for in the training data. This is because the network is able to learn relevant features from the available data, potentially allowing it to generalize and provide accurate estimates even in regions not explicitly covered by the training set. In this sense, the model can adapt to variations in parameters and still maintain robustness in its predictions [11].
3.4 Computation of seismic data
We utilize an in-house forward solver based on the discontinuous Galerkin (DG) method [13] and third-order Adams–Bashforth time-stepping [8] to compute synthetic seismograms. The solver is implemented in C/C++ and parallelized across CPU nodes using the Message Passing Interface (MPI). GPU acceleration is achieved through the OCCA library [24], which enables communication between the CPU and GPU as well as efficient execution on graphics hardware. The computational domain is discretized using tetrahedral elements, and accuracy is controlled via element size and the order of polynomial basis functions. All simulations were performed on the Puhti supercomputing environment maintained by CSC – IT Center for Science, Finland [5], using NVIDIA Volta V100 GPUs. Computational grids used in this work were built using COMSOL Multiphysics.
Two source time functions are used in our modeling: the first derivative of a Gaussian and the Ricker wavelet. Both wavelets are defined below using the same frequency and time delay parameters. We employ two source signatures for the following reason: the first derivative of a Gaussian wavelet is used to generate the training and validation datasets (Section 4.2), whereas the Ricker wavelet is used for the synthetic test set (Section 5.1). This intentional mismatch allows us to assess the ability of the predictive framework to generalize to an unseen source signature.
The first derivative of a Gaussian is defined as
$g=\frac{(t-t_{0})}{c}\exp\left(b\left((t-t_{0})^{2}-c^{2}\right)\right),$ (12)
where $b=-(f\pi)^{2}$ and $c=\sqrt{-0.5/b}$.
The Ricker wavelet is defined as
$g=\left(1+2b(t-t_{0})^{2}\right)\exp\left(b(t-t_{0})^{2}\right),$ (13)
with the same parameters $b$ and $t_{0}$, using a frequency $f=60$ Hz and delay $t_{0}=1.2/f$. Figure 3 (left) shows the source wavelet amplitudes as a function of time.
Figure 3: Left: The two source time functions used in the modeling, first derivative of a Gaussian (solid blue) and Ricker wavelet (dashed red). Right: Cross section of an example mesh used in the simulations.
To ensure accuracy across the simulations, we use polynomial basis functions of order four or five, depending on the dataset, and the mesh resolution is defined relative to the smallest expected wavelength in the model. Specifically, we use a minimum element density criterion of approximately 2.5 elements per shortest wavelength when generating test data (see Section 5.1), while the training and validation sets (see Section 4.2) use slightly lower resolutions, approximately 2.0 and 1.9 elements per wavelength, respectively. This ensures that the datasets have slightly different numerical accuracy, as intended. The purpose of using different discretizations, both in mesh resolution and basis order, is to avoid the so-called inverse crime: if the same forward model and numerical resolution were used for training, validation, and testing, the inversion could become artificially easier and unrealistically accurate. This principle is well established in the general theory of inverse problems [15]. The discretization differences introduce controlled numerical variability. These differences are not intended as noise but ensure that the test data are not generated with the exact same forward operator as the training data. An example mesh used in the synthetic test dataset simulations, shown in Fig. 3 (right), contains 158,180 tetrahedral elements and 28,407 vertices, illustrating the level of discretization applied.